Here's a hard one...HELP
In 1991, th elife expectancy of males in a certain country was 70.4 years. In 1998, it was 74.2 years. Let E represent the life expectancy in year t and let t represent the number of years since 1991.
The linear function E(t) that fits the data is??
E(t) = ??t + ??
(round to the nearest tenth)
Use the function to predict the life expectancy of males in 2003.
E(12)=???
(round to the nearest tenth)

Answer by stanbon(75887)

In 1991, the life expectancy of males in a certain country was 70.4 years.
In 1998, it was 74.2 years.
Let E represent the life expectancy in year t and let t represent the number of years since 1991
-----------
You have two points relating t and E: (0,70.4) and (7,74.2)
slope = (74.2-70.4)/7 = 3.8/7 = 19/35
intercept: 70.4
------------------------
The linear function E(t) that fits the data is??
E(t) =(19/35)t + 70.4
(round to the nearest tenth)
Use the function to predict the life expectancy of males in 2003.
E(12)= (19/35)*12 + 70.4 = 76.9 years
(round to the nearest tenth)
===================
Cheers,
Stan H.